OpenStudy (anonymous):

If two waves are traveling at the same speed, which statement will be true about their wavelengths? A. The wave with higher frequency will have shorter wavelength. B. The wave with lower frequency will have shorter wavelength. C. The wave with higher frequency will have longer wavelength

OpenStudy (amorfide):

You could get the formula and sub in values for example λ=c/f c is the speed, f is the frequency, λ is the wavelength if you have the same speeed c=1 let's say we have one with frequency 1, one with a higher frequency of 2 λ=1/1=1 λ=1/2=0.5 the wave with higher frequency has shorter wavelength

OpenStudy (anonymous):

So its more than likely that it's going to be less wavelength if the frequency of the wave is higher due to the division between speed and frequency ?

OpenStudy (amorfide):

Yes but I am sure there is a more scientific explanation than that somewhere lol